OpenStudy (spring98):

Medals! Simplify (4^7/5^2)^3.

OpenStudy (anonymous):

just plug it into the calculator

OpenStudy (spring98):

sorry my calculator dosn't have all this! can you do it for me and tell me?

OpenStudy (jhannybean):

\[\left(\frac{4^7}{5^2}\right)^3 \iff \frac{(4^7)^3}{(5^2)^3} \iff \frac{4^?}{5^?}\]

OpenStudy (spring98):

4^21/4^6 right? i'm pretty sure thats it!

OpenStudy (dinamix):

4^21/5^6 @Spring98

OpenStudy (spring98):

is the answer?

OpenStudy (jhannybean):

You have a typo in your response @Spring98

OpenStudy (spring98):

4^21/5^6 is the answer? @dinamix

OpenStudy (dinamix):

yup i think so cuz PGCD( 4 , 5 ) = 1 this is answer